# The Value of Information in Human-AI Decision-making

### v4

@@ Abstract
Multiple agents—including humans and AI models—are increasingly combined to make decisions with the expectation of achieving complementary performance, where the decisions they make together outperform those made individually. However, knowing how to improve the performance of collaborating agents is often difficult without knowing more about what particular information and strategies each agent employs. With a focus on human-AI pairings, we contribute a decision-theoretic framework for characterizing the value of information—and consequently, opportunities for agents to better exploit available information–in AI-assisted decision workflows. We present a novel explanation technique (ILIV-SHAP) that adapts SHAP explanations to highlight human-complementing information. We validate the effectiveness of the framework and ILIV-SHAP through a study of human-AI decision-making. We show that our measure of complementary information can be used to identify which AI model will best complement human decisions. We also find that presenting ILIV-SHAP with AI predictions leads to reliably greater reductions in error over non-AI assisted decisions more than vanilla SHAP.  [^1] The code to calculate the main quantities in our framework and reproduce the experimental results is available at https://osf.io/p2qzy/?view_only=bf39de5d96f047f69e45ffd42689ebf9.
@@ 1 Introduction
Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform an AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019b; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). A rigorous understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or agents individually using measures based on post-hoc decision accuracy (Passi and Vorvoreanu, 2022). However, such approaches do not account for the best achievable performance given the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024). Moreover, existing approaches cannot provide insight on the potential for available information in the environment to improve decision performance. Consequently, it is difficult to design interventions that improve the team’s performance.
@@
If we could identify information complementarities that can be exploited to maximize the decision performance of a human and AI model—such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated contextually-available information in their judgments—we can intervene to improve decision-making. For example, if we can identify how much complementary information over human judgments an AI model provides, we can use this knowledge to guide model selection or motivate further data collection to improve the model. Conversely, finding evidence that model predictions contain decision-relevant information that humans do not exploit can motivate the design of explanations highlighting complementary information.
@@
We contribute a decision-theoretic framework for characterizing the value of information available within an AI-assisted decision workflow. In our framework, information is considered valuable to a decision-maker to the extent that it is possible, in theory, to incorporate it into their decisions to improve performance. Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions. While we focus on combinations of humans and AI agents, the methods we propose can be applied to any combination of agents.
@@
We introduce two metrics for evaluating information value in human-AI collaboration. The first—global human-complementary information value—calculates the value of a new piece of information to an agent over all of its possible realizations among all instances. The second—instance-level human-complementary information value—identifies opportunities for decision-makers to better use instance-level information such as specific AI predictions. We derive a new explanation technique (ILIV-SHAP) that reveals how data features influence the value of complementary information for an individual prediction.
@@
To evaluate these tools, we contribute the results of a crowdsourced between-subjects experiment in which humans make decisions with and without AI models with varying human-complementary information and with access to different explanation approaches. We find that an AI model with more human-complementary information leads to greater improvements in human-AI team performance over the human-alone baseline. We also find that adding an ILIV-SHAP explanation to a traditional SHAP explanation leads to greater improvements in human-AI performance over the human-alone baseline than only the SHAP explanation or no explanation. We demonstrate further uses of the framework for model selection, empirical evaluation of human-AI performance, and explanation design on existing datasets spanning medical diagnosis, human deepfake detection, and recidivism prediction (see Appendices E, D and F). We also include an observational study on the dataset from Vodrahalli et al. (2022) which identified the AI model with more human-complementary information helps human-AI teams achieve greater improvements in performance over the human-alone baseline in Appendix C.
@@ 2 Related work
Many empirical studies of human-AI collaboration focus on AI-assisted human decision-making for legal, ethical, or safety reasons (Bo et al., 2021; Boskemper et al., 2022; Bondi et al., 2022; Schemmer et al., 2022). However, a recent meta-analysis by Vaccaro et al. (2024) finds that, on average, human–AI teams perform worse than the better of the two agents alone. In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b, 2019, a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b). The present work differs from much of this prior work by approaching human-AI complementarity from the perspective of information value and use, including asking whether the human and AI decisions provide additional information that is not used by the other.
@@
Our work contributes methods for evaluating the decisions of human-AI teams (Kleinberg et al., 2015, 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025). Kleinberg et al. (2015) proposed that evaluations of human-AI collaboration should be based on the information that is available at the time of decisions. According to this view, our work defines Bayesian best-attainable-performance benchmarks similar to several prior works (Guo et al., 2024; Wu et al., 2023; Agrawal et al., 2020; Fudenberg et al., 2022). Closest to our work, Guo et al. (2024) model the expected performance of a rational Bayesian agent faced with deciding between the human and AI recommendations as the theoretical upper bound on the expected performance of any human-AI team. This benchmark provides a basis for identifying exploitable information within a decision problem.
@@
Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021). Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule. Alur et al. (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts. Our work shares the motivation of incorporating human expertise, but targets a slightly broader scope by quantifying the information value for all available signals and agent decisions in a specific scenario of human–AI decision tasks.
@@ 3 Framework
Our framework takes as input a decision problem associated with an information model and outputs the value of information of any available signals to any agent, conditioning on the existing information in their decisions within a Bayesian decision theoretic framework. The intuition behind our use of statistical decision theory is that while human decision-makers may deviate from Bayesian rational decision-making, understanding the possible value of a signal requires defining the best attainable decision performance with that signal. Our framework provides two separate functions to quantify the value of information: one globally across the data-generating process, and one locally in a realization drawn from the data-generating process. We also introduce a robust analysis approach to information order, which enables us to compare the agent-complementary information in signals for all possible decision problems.
@@
  • A payoff function $S:\mathbf{D}\times\mathbf{\Omega}\to\mathbb{R}$, used to assess the quality of a decision given a realization of the state. For example, $S(d=0,\omega=0)=0,S(d=0,\omega=1)=-100,S(d=1,\omega=0)=-50,S(d=1,\omega=1)=0$, which punishes the DM for selecting an action that does not match the weather.
In decision problems corresponding to prediction tasks, the decision space is a probabilistic belief over the state space, i.e., $\mathbf{D}=\Delta(\mathbf{\Omega})$. For such problems, a payoff function is said to be a proper scoring rule if the optimal action is to predict the true distribution, i.e., $p=\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}]$. For any decision problem with payoff function $S:\mathbf{D}\times\mathbf{\Omega}\to\mathbb{R}$, there is an equivalent proper scoring rule $\hat{S}:\Delta(\mathbf{\Omega})\times\mathbf{\Omega}\to\mathbb{R}$ defined by choosing the optimal decision under the reported belief. Formally,
$\hat{S}(p,\omega)=S(\arg\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim p}\![{S(d,\omega)}],\omega).$ (1)
Equation 1 shows a reduction from the payoff function $S$ to the proper scoring rule $\hat{S}$, i.e., any decision $p$ under $\hat{S}$ represents a decision $d$ in $S$ that best-responds to the distribution $p$. Therefore, the best-attainable performance defined in the proper scoring rule $\hat{S}$ is equivalent to the best-attainable performance defined in any payoff function $S$ that can be reduced to $\hat{S}$. Throughout our framework, we prefer proper scoring rules over non-proper scoring rules, since the best-attainable performance defined in the former implies the best-attainable performance in the latter.
@@
We cast the information available to a DM as a signal defined within an information model. We use the definition of an information model in Blackwell et al. (1951). The information model can be represented by a data-generating model with a set of signals.
  • Signals. There are $n$ “basic signals” represented as random variables $\Sigma_{1},\ldots,\Sigma_{n}$, from the signal spaces $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$. Basic signals represent information available to a decision-maker, e.g., $\mathbf{\Sigma}_{1}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{2}\in\{0,\ldots,100\}$ for temperature in Celsius, etc. The decision-maker observes a signal, which is a subset of the basic signals, $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$. Specifically, we use $V=\{\Sigma_{j_{1}},\ldots,\Sigma_{j_{k}}\}$ for a signal having $k$ basic signals and denote the signal space as $\mathbf{V}=\mathbf{\Sigma}_{j_{1}}\times\ldots\times\mathbf{\Sigma}_{j_{k}}$. For example, a signal representing a combination of two basic signals $V=\{\Sigma_{1},\Sigma_{2}\}$ observed by the decision-maker might consist of cloudiness $\Sigma_{1}$ and the temperature $\Sigma_{2}$ of the day. Given a signal composed of $m$ basic signals, we write the realization of $V$ as $v=(\sigma_{j_{1}},\dots,\sigma_{j_{k}})$, where the realizations $\sigma_{j_{i}}\in\mathbf{\Sigma}_{j_{i}}$ are sorted by the index of the basic signals $j_{i}\in[n]$. The union $V$ of two signals $V_{1},V_{2}$ takes the set union, i.e., $V=V_{1}\cup V_{2}$.
  • Data-generating process. A data-generating process is a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega})$ over the basic signals and the payoff-relevant state. $\pi$ can be viewed as the combination of two distributions: the prior distribution of the state $\Pr[\omega]$ and the signal-generating distribution $\Pr[v|\omega]$ defining the conditional distribution of signals. The DM may only observe a subset $V$ of the $n$ basic signals. We can define the Bayesian posterior belief upon receiving a signal $V=v$ from the data-generating model as
@@
    where we slightly abuse notation to write $\pi(v,\omega)$ as the marginal probability of the signal realized to be $v$ and the state being $\omega$ with expectation over other signals.
The choice of basic signals directly impacts how many observed samples are required to get a good estimate of the data-generating process. When high-dimensional signals such as images or text are used, it may not be computationally feasible to estimate the data-generating process. In such cases, one can preprocess the high-dimensional signals to get low-dimensional representations which (ideally) capture any important structure. These lower-dimensional signals can be defined by humans, such as when concepts are identified and then used to label the high-dimensional signals (e.g., images or parts of images). Alternatively, they can be defined algorithmically, by strategically applying dimensionality reduction to generate low-dimensional embeddings. We introduce an algorithm for identifying a set of optimal “coarsened” signals given a set of observed samples in Appendix B.
@@
$\mathrm{R}^{\pi,S}(V)=\text{\bf E}_{(v,\omega)\sim\pi}\![{S(d^{r}(v),\omega)}]$ (2)
@@
We use $\emptyset$ to represent a null signal, such that $\mathrm{R}^{\pi,S}(\emptyset)$ is the expected payoff of a Bayesian rational DM who only uses the prior distribution to make decisions. In this case, the Bayesian rational DM will take the best fixed action under the prior, and their expected payoff is:
$\mathrm{R}^{\pi,S}(\emptyset)=\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi}\![{S(d,\omega)}]$ (4)
This baseline defines the maximum expected payoff that can be achieved with no information. Bayesian decision theory quantifies the information value of $V$ as the payoff improvement of $V$ over the payoff obtained without information.
@@
$\operatorname{IV}^{\pi,S}(V)=\mathrm{R}^{\pi,S}(V)-\mathrm{R}^{\pi,S}(\emptyset)$ (5)
We adopt the same idea to define the agent-complementary information value in our framework.
@@ 3.1 Agent-Complementary Information Value
With the above definitions, it is possible to measure the best case additional value that new signals can provide over the information already captured by an agent’s decisions. Here, agent may refer to a human, an AI system, or a human–AI team. The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their decisions. We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM. We model the agent decisions as a random variable $D^{b}$ from the action space $\mathbf{D}$, which follows a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega}\times\mathbf{D})$ with the state and signals. The expected payoff of a Bayesian rational DM who knows $\pi$ is given by the function:
$\mathrm{R}^{\pi,S}(D^{b})=\text{\bf E}_{(d^{b},\omega)\sim\pi}\![{S(d^{r}(d^{b}),\omega)}]$
We seek to identify signals $V$ that can potentially improve agent decisions by analyzing the information value in the combined signal $D^{b}\cup V$ over the information value in $D^{b}$, which we define as the agent-complementary information value.
@@
$\operatorname{ACIV}^{\pi,S}(V;D^{b})=\mathrm{R}^{\pi,S}(D^{b}\cup V)-\mathrm{R}^{\pi,S}(D^{b})$ (6)
@@
Furthermore, $\operatorname{ACIV}$ can reveal complementary information between different types of agents. For instance, if we view AI predictions as $V$ and treat human decisions as the agent signal $D^{b}$, a large $\operatorname{ACIV}$ indicates that AI predictions add considerable value beyond what humans alone achieve. In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans can contribute over the information captured in the AI predictions. We demonstrate uses of $\operatorname{ACIV}$ in Appendix D and Appendix E.

### v5

@@ Abstract
Multiple agents are increasingly combined to make decisions with the expectation of achieving complementary performance, where the decisions they make together outperform those made individually. However, knowing how to improve the performance of collaborating agents requires knowing what information and strategies each agent employs. With a focus on human-AI pairings, we contribute a decision-theoretic framework for characterizing the value of information. By defining complementary information, our approach identifies opportunities for agents to better exploit available information in AI-assisted decision workflows. We present a novel explanation technique (ILIV-SHAP) that adapts SHAP explanations to highlight human-complementing information. We validate the effectiveness of ACIV and ILIV-SHAP through a study of human-AI decision-making, and demonstrate the framework on examples from chest X-ray diagnosis and deepfake detection. We find that presenting ILIV-SHAP with AI predictions leads to reliably greater reductions in error over non-AI assisted decisions more than vanilla SHAP [^1] The code to calculate the main quantities in our framework and reproduce the experimental results is available at https://osf.io/p2qzy/?view_only=bf39de5d96f047f69e45ffd42689ebf9..
@@ 1 Introduction
Evidence of complementary performance between humans and AI nonetheless remains limited, with many studies showing that human-AI teams underperform the AI alone (Buçinca et al., 2020; Bussone et al., 2015; Green and Chen, 2019; Jacobs et al., 2021; Lai and Tan, 2019; Vaccaro and Waldo, 2019; Kononenko, 2001). A rigorous understanding of such results is limited by the fact that most analyses of human-AI decision-making focus on ranking the performance of human-AI teams or agents individually using measures based on post-hoc decision accuracy (Passi and Vorvoreanu, 2022). However, such approaches do not account for the best achievable performance given the information available at the time of the decision (Kleinberg et al., 2015; Guo et al., 2024; Rambachan, 2024). Moreover, existing approaches cannot provide insight on the potential for available information in the environment to improve performance. Consequently, it is difficult to design interventions that improve the team’s performance.
@@
Imagine one could identify information complementarities that can be exploited, such as when one of the agents has access to information not contained in the other’s judgments, or has not fully integrated contextually-available information in their judgments. This would motivate interventions to improve decision-making. For example, if we can identify how much complementary information AI predictions provide over human judgments, we can use this knowledge to guide model selection or motivate further data collection to improve the model. Conversely, finding evidence that model predictions contain decision-relevant information that humans do not exploit can motivate the design of explanations communicating complementary information.
@@
We contribute a decision-theoretic framework for characterizing the value of information available in an AI-assisted decision workflow. In our framework, information is considered valuable to a decision-maker to the extent that it is possible, in theory, to incorporate it into their decisions to improve performance. Specifically, our approach analyzes the expected marginal payoff gain from best case (Bayes rational) use of additional information over best case use of the information already encoded in agent decisions. The rational framework allows us to upperbound the expected payoff that is achievable by any strategies in the same experiment, and identifies sub-optimality in agent use of information by comparing to rational behavior. The upper bound our framework estimates holds regardless of the human’s decision-making process, which may deviate from rationality, given a specified decision problem. Further, our methods can be used even when the decision problem definition is ambiguous, by using a robustness analysis over all possible proper scoring rules to identify the upper bound of performance in the worst case.
@@
We introduce two metrics for evaluating information value in human-AI collaboration. The first—global human-complementary information value—calculates the value of a new piece of information to an agent over all of its possible realizations among all instances. The second—instance-level human-complementary information value—identifies opportunities for decision-makers to better use instance-level information such as specific AI predictions. Applying the second metric, we derive a new explanation technique (ILIV-SHAP) that reveals how data features influence the value of complementary information for an individual prediction.
@@
To evaluate these tools, we contribute the results of a crowdsourced between-subjects experiment in which humans make decisions with and without AI models with varying human-complementary information and different explanation approaches. We find that an AI model with more human-complementary information leads to greater improvements in human-AI team performance over the human-alone baseline. We also find that adding an ILIV-SHAP explanation to a traditional SHAP explanation leads to greater improvements in human-AI performance over the human-alone baseline than only the SHAP explanation or no explanation. We demonstrate further uses of the framework in real-world decision-making tasks, including chest X-ray diagnosis (Rajpurkar et al., 2018; Johnson et al., 2019) and deepfake detection (Dolhansky et al., 2020; Groh et al., 2022) [^2] We also include an observational study on the dataset from Vodrahalli et al. (2022) which identified the AI model with more human-complementary information helps human-AI teams achieve greater improvements in performance over the human-alone baseline. See Appendix C..
@@ 2 Related work
Many empirical studies of human-AI collaboration focus on AI-assisted human decision-making for legal, ethical, or safety reasons (Bo et al., 2021; Boskemper et al., 2022; Bondi et al., 2022; Schemmer et al., 2022). However, a recent meta-analysis by Vaccaro et al. (2024) finds that, on average, human–AI teams perform worse than the better of the two agents alone. In response, a growing body of work seeks to evaluate and enhance complementarity in human–AI systems (Bansal et al., 2021b, 2019, a; Wilder et al., 2021; Rastogi et al., 2023; Mozannar et al., 2024b). The present work differs from much of these prior works by approaching human-AI complementarity from the perspective of information value and use, including asking whether the human and AI decisions provide additional information that is not used by the other.
@@
Our work contributes methods for evaluating the decisions of human-AI teams (Kleinberg et al., 2015, 2018; Lakkaraju et al., 2017; Mullainathan and Obermeyer, 2022; Rambachan, 2024; Guo et al., 2024; Ben-Michael et al., 2024; Shreekumar, 2025). Kleinberg et al. (2015) proposed that evaluations of human-AI collaboration should be based on the information that is available at the time of the decision. According to this view, our work defines Bayesian best-attainable-performance benchmarks similar to several prior works (Guo et al., 2024; Wu et al., 2023; Agrawal et al., 2020; Fudenberg et al., 2022). Closest to our work, Guo et al. (2024) model the expected performance of a rational Bayesian agent faced with deciding between the human and AI recommendations as the theoretical upper bound on the expected performance of any human-AI team. This benchmark provides a basis for identifying exploitable information within a decision problem.
@@
Some approaches focus on automating the decision pipeline by explicitly incorporating human expertise in developing machine learning models, such as by learning to defer (Mozannar et al., 2024a; Madras et al., 2018; Raghu et al., 2019; Keswani et al., 2022, 2021; Okati et al., 2021; Chen et al., 2022). Corvelo Benz and Rodriguez (2023) propose multicalibration over human and AI model confidence information to guarantee the existence of an optimal monotonic decision rule. Alur et al. (2023) propose a hypothesis testing framework to evaluate the added value of human expertise over AI forecasts. Our work shares the motivation of incorporating human expertise, but provides a general analytical framework for quantifying the information value of all available signals and agent decisions in human–AI decision tasks.
@@ 3 Framework
Our framework takes as input a decision problem associated with an information model, including decisions from one or more agents. It outputs the value of information of available signals to the agents, conditioning on the existing information in their decisions. Understanding the possible value of a signal to a decision-maker requires defining the best attainable decision performance with that signal. We therefore adopt a Bayesian decision theoretic framework.
Our framework provides two separate functions to quantify the value of information: one globally across the data-generating process, and one locally in a realization.
@@
  • A payoff function $S:\mathbf{D}\times\mathbf{\Omega}\to\mathbb{R}$, used to assess the quality of a decision given a realization of the state. For example, $S(d=0,\omega=0)=0,S(d=0,\omega=1)=-100,S(d=1,\omega=0)=-50,S(d=1,\omega=1)=0$, which punishes the DM for wrongly taking or not taking the umbrella.
@@
We cast the information available to a DM,including any available agent decisions, as a set of signals defined within an information model. Following the definition of an information model in Blackwell et al. (1951), the information model can be represented by a data-generating model with a set of signals.
  • Signals. There are $n$ “basic signals” represented as random variables $\Sigma_{1},\ldots,\Sigma_{n}$, from the signal spaces $\mathbf{\Sigma}_{1},\ldots,\mathbf{\Sigma}_{n}$. Basic signals represent information available to a decision-maker as they decide, e.g., $\mathbf{\Sigma}_{1}\in\Delta\mathbf{\Omega}$ for a probabilistic prediction of raining, $\mathbf{\Sigma}_{2}=\{\text{cloudy},\text{not cloudy}\}$, $\mathbf{\Sigma}_{3}\in\{0,\ldots,100\}$ for temperature in Celsius, etc. The decision-maker observes a signal, which is a subset of the basic signals, $V\subseteq 2^{\{\Sigma_{1},\dots,\Sigma_{n}\}}$. The combination of two signals $V_{1}$ and $V_{2}$ takes the set union $V=V_{1}\cup V_{2}$. For example, in the standard human-AI decision workflow where a human makes an independent judgment before consulting the AI, the set of all basic signals is $V=\{x,D^{\text{H}},D^{\text{AI}},e(D^{\text{AI}},x)\}$–the features of the instance, the human’s initial decisions, the AI’s predictions, and any explanation of the AI’s predictions.
  • Data-generating process. Decision benchmarks are defined relative to a specific data-generating process, a joint distribution $\pi\in\Delta(\mathbf{\Sigma}_{1}\times\ldots\times\mathbf{\Sigma}_{n}\times\mathbf{\Omega})$ over the basic signals and the payoff-relevant state. $\pi$ can be viewed as the combination of two distributions: the prior distribution of the state $\Pr[\omega]$ and the signal-generating distribution $\Pr[v|\omega]$ defining the conditional distribution of signals. To represent the best-attainable performance of observing a subset $V$ of the $n$ basic signals, we use the Bayesian posterior belief upon receiving a signal $V=v$ as
@@
    where $\pi(v,\omega)$ denotes the marginal probability of the signal realized to be $v$ and the state being $\omega$ with expectation over other signals, and $\pi(\omega)$ denotes the prior $\Pr[\omega]$. Throughout the paper, we use capital letters to denote a random variable of signal, e.g., $V$, and use little letetrs to denote a realization of signal, e.g., $v$.
@@
$\mathrm{R}(V):=\text{\bf E}_{(v,\omega)\sim\pi}\![{S(d^{r}(v),\omega)}]$ (1)
@@
We further characterize the maximum expected payoff that can be achieved with no information. This can be used as the baseline to quantify the information value of $V$ as the payoff improvement of $V$ over. We use $\emptyset$ to represent a null signal and $\mathrm{R}(\emptyset)$ represents the expected payoff of a Bayesian rational DM who only uses the prior distribution to make decisions. In this case, the Bayesian rational DM takes the best fixed action under the prior, and their expected payoff is:
$\mathrm{R}(\emptyset):=\max_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\pi}\![{S(d,\omega)}]$ (3)
@@
$\operatorname{IV}(V):=\mathrm{R}(V)-\mathrm{R}(\emptyset)$ (4)
We adopt the same idea below to define the agent-complementary information value.
@@ 3.1 Agent-Complementary Information Value
With the above definitions, it is possible to measure the best case additional value that new signals can provide over the information already captured by an agent’s decisions. Here, agent may refer to a human, an AI system, or a human–AI team. The intuition behind our approach is that any information that is used by decision-makers should eventually reveal itself through variation in their decisions. We recover the information value in agent decisions by offering the decisions as a signal to the Bayesian rational DM.
@@
$\operatorname{ACIV}(V;D^{b}):=\mathrm{R}(D^{b}\cup V)-\mathrm{R}(D^{b})$ (5)
@@
Furthermore, $\operatorname{ACIV}$ can reveal complementary information between different types of agents. For instance, if we view AI predictions as $V$ and treat available human decisions as the agent signal $D^{b}$, a large $\operatorname{ACIV}$ indicates that AI predictions add considerable value beyond what humans alone achieve. In the reverse scenario, if human decisions serve as $V$ and AI predictions are $D^{b}$, we can measure how much humans can contribute over the information captured in the AI predictions.
Of course, we are unlikely to observe identical realizations of the signals in continuous-valued or high-dimensional data, such as images or text. A natural relaxation is to consider the realizations that are sufficiently “similar” for the performance of the decision-maker. Algorithm 1 learns the posterior distribution of the payoff-relevant state from the signal and then uses that prediction as the probability distribution of the payoff-relevant state to choose the optimal action. This approach can be readily extended to the case where the signal is more complex, such as human decisions in the form of freeform text (e.g., radiology reports) or explanations of the AI predictions in the form of images (e.g., saliency map).
Algorithm 1 A method to calculate the ACIV of $V$.
1:Input: Observed realizations $\{v_{i},d^{b}_{i},\omega_{i}\}_{i=1}^{n}$, a regression algorithm $\mathcal{A}$, payoff function $S$, decision space $\mathbf{D}$, and state space $\mathbf{\Omega}$.
2:Output: $\operatorname{ACIV}(V;D^{b})$
[…]
5:for $i=1$ to $n$ do
6:  $\hat{p}_{i}\leftarrow\hat{a}(v_{i},d^{b}_{i})$
7:  $\hat{d}^{r}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}_{i}}\![{S(d,\omega)}]$
8:  $\hat{p}^{b}_{i}\leftarrow\hat{a}^{b}(d^{b}_{i})$
9:  $\hat{d}^{rb}_{i}\leftarrow\operatorname{argmax}_{d\in\mathbf{D}}\text{\bf E}_{\omega\sim\hat{p}^{b}_{i}}\![{S(d,\omega)}]$
10:  $s_{i}\leftarrow S(\hat{d}^{r}_{i},\omega_{i})-S(\hat{d}^{rb}_{i},\omega_{i})$
11:end for
12:return $\frac{1}{n}\sum_{i=1}^{n}s_{i}$